A steel piano wire, of length 0.750 m and mass 4.40 g is stretched under a tension of 630.0 N. What is the speed of transverse waves on the wire?
Answer #1

length of wire = 0.75

mass = 4.4 g

mass per unit length u= m/ L =4.4 /0.75= 5.867 g/m

tension in the wire = 630 N

v = sqrt(T/u)

= sqrt(630 / 0.005867)

= 327.7 m/s

so the speed of the transverse wave is 327.7 m/s